# Pairing in Fission: Studies of Static Approach and Collective Inertias

### v1

@@ Abstract
Pairing plays a crucial role in the microscopic description of nuclear fission. Microscopic methods provide access to three quantities related to pairing, namely, the pairing gap ($\Delta$), the particle number fluctuations ($\Delta\hat{N}^{2}$), and the quenching factor (QF). The aim of this work is to analyse the impact of each of these quantities on the static description of the fission process.
@@ I Introduction
The nuclear fission process discovered more than 80 years ago Hahn and Straßmann (1939) still remains one of the most fascinating and not fully understood nuclear phenomenon Schmidt and Jurado (2018); Bender et al. (2020). Shortly after the first observation of fission in the laboratory, the process was given an explaination based on the liquid drop model assumption of the nucleus Meitner and Frisch (1939) that is still valid nowadays. The subsequent discovery one year later of spontaneous fission Flerov and Petrzhak (1940), indicated that the process itself is much more complex than initially thought. These early studies opened up the door to a myriad of fascinating experimental and theoretical studies that have been carried out and published since then.
@@
The non-negligible role of pairing correlations in fission dynamics strongly affects the calculation of half-lives. Two main factors play a crucial role in this relationship. First, modification of the fission barrier height and nuclear deformation at the barrier. Second, reduction of the collective inertia during the increase of deformation from ground state to scission. These aspects have been extensively studied in the literature, employing a large variety of nuclear models.
@@
The first research on the impact of pairing interaction on fission was conducted in the 1970s when Ledergerber and Pauli Ledergerber and Pauli (1973) determined fission paths using the liquid drop model with Strutinski shell correction and Bardeen-Cooper-Schrieffer (BCS) surface pairing in the space of deformation, including asymmetric shapes. The authors showed that the value of pairing gap $\Delta$ in the BCS model strongly affects the inertia parameter. In consequence, pairing interaction leads to the reduction of the penetration integral at large deformation without modifying the potential energy.
@@
Moreto and Babinet  Moretto and Babinet (1974) applied a simple, realistic model to show that an increase in the pairing gap parameter leads to an increase in the penetrability of the fission barrier. The main message from this study is that there is a strong interplay between the potential energy, which increases as a square of the pairing gap, and the collective inertia, which decreases as an inverse of the pairing gap squared. As a consequence, the least action path leads through regions with a larger pairing gap than the least energy path.
@@
The influence of pairing interaction on fission observables has been continued in a series of papers of the group from Maria Curie-Sklodowska University in Lublin Staszczak et al. (1985); Łojewski and Pomorski (1980); Staszczak et al. (1989); Łojewski and Staszczak (1999). The fission half-lives have been calculated in various macroscopic-microscopic approaches with BCS pairing. Proton ($\Delta_{p}$) and neutron ($\Delta_{n}$) pairing gap parameters were treated on the same footing as deformation parameters. The dynamical fission path was calculated by minimizing the action integral in the multidimensional space that included pairing degrees of freedom Staszczak et al. (1985); Łojewski and Staszczak (1999). It was shown that in 252Fm, the shortest half-lives are obtained for the fission path with larger $\Delta$ than in the static path that minimizes energy. Despite the increased size of the fission barrier by 1.4 MeV, the mass parameter along the dynamic path was reduced by half, leading to the reduction of half-lives by three orders of magnitude. A similar effect was observed along the fermium isotopic chain.
@@
The impact of the quadrupole pairing on the spontaneous fission half-lives of Fm isotopes has been studied in Łojewski and Pomorski (1980) in the static approach. The authors observed the decrease of the second barrier height by 1 MeV for the quadrupole pairing forces, whereas the mass parameters may increase, especially on the first fission barrier. In some Fm isotopes, including quadrupole together with the monopole pairing leads to a reduction of half-lives by two orders of magnitude.
@@
In Ref. Staszczak et al. (1989), the generator coordinate method, together with the Gaussian overlap approximation (GCM+GOA), has been applied to investigate the influence of pairing vibrations on the spontaneous fission half-lives calculations of even-even transplutonium isotopes. The authors found that the action integrals calculated along the static fission paths are around 13% larger than the integrals obtained from dynamical trajectories. Again, the reduction of spontaneous fission half-lives is the result of the decrease of mass parameters, which compensates the increase of fission barriers. The dynamic paring allowed to reduce the calculated half-lives by several orders of magnitude.
The findings mentioned above from macroscopic-microscopic models have also been observed in self-consistent studies. The authors of Ref. Karatzikos et al. (2010) presented the study of spontaneous fission barriers of heavy and super-heavy nuclei within the RHB theory and Relativistic Mean Field (RMF) plus BCS methods. The authors obtained an increase in the absolute value of the pairing energy and a strong decrease in barrier height with the increase of the pairing $\Delta$ along the fission path relative to its ground-state value.
@@
In Ref. Sadhukhan et al. (2014); Giuliani et al. (2014); Zhao et al. (2016); Rodríguez-Guzmán et al. (2023), the impact of pairing on fission paths was studied using the particle-number dispersion $\Delta\hat{N}^{2}$ as collective degree of freedom. These studies, which involved Skyrme, Gogny, and relativistic energy density functionals, showed that the inclusion of $\Delta\hat{N}^{2}$ strongly modifies the least-action paths, producing an overall reduction of spontaneous fission lifetimes compared to static calculations and dynamical calculations where only mass multipoles are considered as collective degrees of freedom. Additionally, it has been demonstrated that in some nuclei pairing as a dynamic degree of freedom restores the axial symmetry at the first fission barrier Sadhukhan et al. (2014); Zhao et al. (2016). Time-dependent GCM+GOA calculation employing $\Delta\hat{N}^{2}$ as collective degree of freedom have shown that pairing correlations play an important role in shaping the fission fragment yields in induced fission Zhao et al. (2021). Also, $\Delta\hat{N}^{2}$ has been employed in beyond-mean field studies exploring the role of dynamic pairing correlations in fission Rodríguez-Guzmán et al. (2023).
An alternative approach to controlling the pairing interactions has been proposed in the Refs. Kouno et al. (2022, 2021) The method for modification of the strength of pairing interaction was achieved by changing the $G$ parameter in the RMF+BCS calculations. It was shown that increasing the pairing strength up to 20% reduces the fission barrier by around 2.5 MeV in Pu isotopes due to the higher single-particle level densities around the Fermi surface at the saddle point. This result is in contradiction to the research presented above when dynamic fission paths are characterized by stronger pairing and higher barriers at the same time. Recently, this method was intensively used in the analysis of the pairing interaction during the fission.
A similar relationship was found in Refs. Guan et al. (2021, 2023), where the authors employed a macroscopic-microscopic model with BCS pairing. It has also been found that the ground states and the scission configurations are less sensitive to the pairing strength than the saddle points. These results were confirmed in energy density functional studies employing the Skyrme forces in the fission of 240Pu Chen et al. (2022). The authors observed a decrease of the fission barriers up to 2.3 MeV in both symmetric and asymmetric fission channels when the pairing strength is increased by 10%. The deformation of a nucleus at the scission line can also be slightly modified by variation of pairing strength. A similar result has been presented in Ref.  Wang et al. (2023).
In microscopic Hartree-Fock-Bogoliubov (HFB) calculations, the impact of pairing correlations has also been studied by modifying the strength of the pairing field Giuliani and Robledo (2013); Rodríguez-Guzmán and Robledo (2014a, b); Guzman and Robledo (2017). These studies have shown that increasing the pairing strength by 5%-10% can reduce the spontaneous fission lifetimes by several orders of magnitudes, particularly for those nuclei with broad and high fission barriers, mostly because of the strong dumping of the collective inertias. Moreover, when pairing is stronger, the mass parameters show smoother behavior without sharp peaks along the fission path.
@@
As the literature review presented here shows, in the last fifty years, the role of pairing in fission has been extensively studied. The non-negligible impact of pairing in fission barriers and half-lives is currently well known. However, as various theoretical methods were used to modify the strength of the pairing interaction, results obtained from different studies cannot always be directly compared. Moreover, some contradictory conclusions were drawn from these studies, namely differences in the modification of the size of the fission barrier with pairing. These ambiguities indicate the need to systematize the knowledge regarding the impact of pairing in static fission calculations. In this work, we closely examine the three self-consistent model parameters corresponding to pairing and thoroughly investigate their impact on the description of the fission process, employing a consistent theoretical framework based on the HFB calculations. The paper is organized as follows: Section II briefly presents the theoretical model employed in this work. Section III presents the results of our studies. We first address the impact of the pairing gap, the particle number fluctuations, and the quenching factor parameters on fission barriers. We then discuss the results of fission half-lives for various pairing strengths and different aspects of collective inertia parameters along fission paths. In Section IV, we summarize our main conclusions. In Appendix A, we discuss the problem of the relative sign of the occupancies in the Bogoliubov transformation.
@@ II Theoretical framework
The starting point for our work is the computation of potential energy surface (PES). We use the microscopic, self-consistent HFB model with the finite-range Gogny-type nuclear interaction. We applied the widely used and well-established D1S parametrization Gogny (1973, 1975); Dechargé and Gogny (1980); Berger et al. (1989, 1991). The computations were done in a one-center axially symmetric, deformed harmonic oscillator basis with 16 shells (including as many as 23 one-dimensional harmonic oscillator shells along the symmetry axis $n_{z\,\textup{max}}=22$). The spontaneous fission half-lives are computed using the Wentzel-Kramers-Brillouin (WKB) approximation, assuming the axially symmetric quadrupole moment $Q_{20}$ as the leading coordinate. The estimation of the probability of tunneling through the one-dimensional fission barrier requires the computation of the action $S$ along the least energy path Brack et al. (1972):
@@
being in and out the classical turning points defined by $V(Q_{20})=E_{0}$. The potential energy in the expression of the action $V(Q_{20})$ is the sum of the HFB energy and the rotational and zero point energy correction Egido and Robledo (2004); Ring and Schuck (1980). The $E_{0}$ parameter represents the quantal ground state energy obtained by solving the collective Schrödinger equation with the quadrupole moment as a degree of freedom. Although this parameter can be easily estimated by assuming a parabolic behavior of the PES around the ground state minimum and considering the collective inertia at that point, we employ a more pragmatic approach and take it equal to 0.5 MeV. $B(Q_{20})$ is the collective inertia calculated within the perturbative cranking approximation Baran et al. (2011):
@@
where $M_{-n}$ are the so-called moments of order $n$ of the collective coordinate $Q_{20}$ (see Baran et al. (2011) for the definition). The formula for the inertia is based on the assumption that the collective operator is a one-body one, which is not true for the fluctuation on the number of particles operator $\Delta\hat{N}^{2}$. To overcome this problem, the traditional approach is to take the one-body part in the expression of $\Delta\hat{N}^{2}$ in the quasiparticle basis. Although this seems a reasonable assumption, it is desirable to test it against the results obtained from a truly one-body operator, like the pairing gap one discussed below. Within the WKB approximation, the spontaneous fission half-life (in seconds) is given by Schunck and Robledo (2016):
@@
As mentioned in the Introduction I, the goal of the present work is to discuss the impact of pairing correlations on the quantities entering the action integral (2). To do so, our analysis covers the three main methods that have been traditionally employed to characterize pairing correlations strengths in fission: HFB calculations with constraints on the pairing gap parameter ($\Delta$), on the particle number fluctuation operator ($\Delta\hat{N}^{2}$), and the renormalization of the pairing interaction.
The pairing gap operator can be defined as $\Delta=\frac{G}{2}(\hat{P}+\hat{P}^{+})$, being $\hat{P}^{+}$ the Cooper pair creation operator, and $G=10/A$ (MeV), a parameter related to the typical strength of the monopole pairing interaction. By employing $\Delta$ as constraining operator, it is thus possible to control the amount of pairing correlations in HFB wave functions $|\Phi\rangle$. The main advantage of employing $\Delta$ a collective degree of freedom is its one-body nature, which allows a consistent estimation of collective inertias through Eq. (2). However, in order to employ $\Delta$ as a collective degree of freedom, specific numerical strategies must be followed in order to properly evaluate the expectation value, as detailed in the Appendix A.
As one of the consequences of the existence of pairing correlations in a HFB calculation is the spontaneous symmetry breaking of the particle number symmetry, it is possible to employ one of the associated order parameters as a measure of pairing correlations. The most straightforward order parameter is associated with the particle number fluctuation $\Delta\hat{N}^{2}=(\hat{N}-\langle\hat{N}\rangle)^{2}$, being $\hat{N}$ the particles number operator. This quantity is zero in the absence of pairing correlations and becomes relatively large (of the order of a few units) for strongly correlated systems. As mentioned before, a drawback related to the fluctuation in particle number is its two-body-operator character. However, owing to the gradient method used in the solutions of the HFB equation, two-body constraint operators can be handled on the same foot as one-body ones.
As a last measure of pairing correlations, we consider the so-called quenching factor (QF), a parameter that renormalizes the strength of the pairing interaction of the Gogny force. The QF is introduced by multiplying the HFB pairing field $\Delta^{(\tau)}$ ($\tau=p,n$) by a constant factor in each iteration of the iterative solution of the HFB equation. This QF parameter is equal to 1.0 for not-modified HFB fields and takes typical values around 1.1 for enhanced pairing. It can be pictured as an analog of the pairing strength $G$ of other simple models, like BCS. In calculations employing the QF,
pairing correlations can be characterized by the so-called particle-particle energy $E_{pp}$ for protons and neutrons, which in HFB formalism is given by:

### v2

# Pairing in Fission: Mean-Field and Collective Inertias Study

@@ Abstract
Pairing plays a crucial role in the microscopic description of nuclear fission. Microscopic methods provide access to three quantities related to pairing, namely, the pairing gap ($\Delta$), the particle number fluctuations ($\Delta\hat{N}^{2}$), and the quenching factor (QF) of the pairing strength. The aim of this work is to analyse the impact of each of these quantities on the mean-field description of the fission process, including collective inertias.
@@ I Introduction
In spite of having been discovered more than 80 years ago  Hahn and Straßmann (1939), the nuclear fission process still remains one of the most fascinating and not fully understood nuclear phenomenon Schmidt and Jurado (2018); Bender et al. (2020). Shortly after the first observation of fission in the laboratory, the process was explained using the liquid drop model assumption of the nucleus Meitner and Frisch (1939) and this interpretation is still valid nowadays. The subsequent discovery of spontaneous fission Flerov and Petrzhak (1940), indicated that the fission process itself is much more complex than initially thought. These initial studies opened up the door to a myriad of fascinating experimental and theoretical studies that have been carried out, and published since then.
@@
It is widely recognized that the non-negligible role of pairing correlations in fission dynamics strongly affects the calculation of half-lives. Two main factors play a crucial role in this connection. First, the modification of the fission barrier height and nuclear deformation at the barrier induced by pairing. Second, the reduction of the collective inertia along the path connecting the ground state with the scission point. These two aspects have been extensively studied in the literature, employing a large variety of nuclear models and interactions.
@@
The first study on the impact of pairing correlations on fission was conducted in the early 1970s. Brack et al. noticed that mass parameter in fission process follows almost exactly a $1/\Delta^{2}$ law (here $\Delta$ is the pairing gap) due to the impact on level density near the Fermi energy induced by pairing Brack et al. (1972). They realized that fission half-lives may be significantly lowered when the fission path does not follow the lowest energy path in the potential energy, and it does not necessarily pass through the extremal points of the deformation energy. Soon after, Ledergerber and Pauli Ledergerber and Pauli (1973) determined fission paths using the liquid drop model with Strutinski shell correction and (surface) pairing treated in the Bardeen-Cooper-Schrieffer (BCS) approach. The authors noticed the strong dependence of the inertia parameter with respect to the value of pairing gap $\Delta$ in the BCS model. In consequence, the consideration of pairing leads to the reduction of the tunneling probability through the fission barrier without modifying the potential energy.
@@
Moreto and Babinet  Moretto and Babinet (1974) applied a simple, realistic model to show that an increase in the pairing gap parameter leads to an increase in the penetrability of the fission barrier. The main message from this study is that there is a strong interplay between the potential energy, which increases as a square of the pairing gap, and the collective inertia, which decreases as an inverse of the pairing gap squared. As a consequence, the least-action path leads through regions with a larger pairing gap than the least-energy path.
@@
The study of the influence of pairing on fission observables was extensively explored in a series of papers coming from the nuclear physics group at Maria Curie-Sklodowska University in Lublin Staszczak et al. (1985); Łojewski and Pomorski (1980); Staszczak et al. (1989); Łojewski and Staszczak (1999). The fission half-lives were computed using various macroscopic-microscopic approaches, including pairing treated in the BCS framework. Proton ($\Delta_{p}$) and neutron ($\Delta_{n}$) pairing gap parameters were treated on equal footing as the standard shape deformation parameters. Two types of fission paths were considered: static (also known as least-energy) path as a result of the minimization of the energy on the way from ground state to scission and dynamic (also named least-action) path where the action integral along the path is minimized. The second one is more demanding computationally as it requires simultaneous analysis of both energy and collective inertia at each point of the space of (deformation) parameters. In the mentioned papers, the dynamical fission path was determined by minimizing the action integral in the multidimensional space that included pairing degrees of freedom Staszczak et al. (1985); Łojewski and Staszczak (1999). This is because of the strong dependence of the inertia with the pairing gap discussed above. As an example, it was shown that in 252Fm, the shortest half-lives are obtained for the fission path with $\Delta$ values larger than those in the static path minimizing the energy. In the dynamic path the height of the fission barrier increases by 1.4 MeV with respect to the result for the static path but the mass parameter along the dynamic path gets reduced by a factor of two, leading to a reduction of half-lives of three orders of magnitude. A similar effect was observed for other members of the fermium isotopic chain.
@@
In the previous papers, only the monopole component of pairing was considered. The impact of quadrupole pairing on the spontaneous fission half-lives of Fm isotopes was studied in Łojewski and Pomorski (1980) in the static approach. Quadrupole pairing decreases the height of the second barrier by 1 MeV, whereas the mass parameter tends to increase, especially on the first fission barrier. Overall, including quadrupole pairing leads to a reduction in the half-lives of up to two orders of magnitude in some of the Fm isotopes considered.
@@
In Ref. Staszczak et al. (1989), the generator coordinate method, together with the Gaussian overlap approximation (GCM+GOA), was applied to investigate the influence of pairing vibrations on the spontaneous fission half-lives calculations of even-even transplutonium isotopes. The authors found that the action integrals calculated along the static fission paths are around 13% larger than the integrals obtained from dynamical trajectories. This is a consequence of the decrease of mass parameters, which compensates the increase of fission barriers. Considering dynamical fission paths with pairing as a collective variable reduces the calculated half-lives by several orders of magnitude. The authors of Ref. Pomorski and Ivanyuk (2009) noticed that deformation dependent pairing energy decreases barrier heights in light nuclei improving agreement with experimental data.
Previous findings are also observed in studies using self-consistent mean field methods. In the framework of the Relativistic Hartree-Bogoliubov (RHB) and Relativistic Mean Field (RMF) plus BCS methods it was shown in Ref. Karatzikos et al. (2010) the strong dependence of pairing gaps with deformation (invalidating the assumption of using a constant gap) as well as the impact on the barrier heights depending on the use of zero or finite range pairing forces.
@@
In Refs. Giuliani et al. (2014); Sadhukhan et al. (2014); Zhao et al. (2016); Rodríguez-Guzmán et al. (2023), the impact of pairing on fission was studied using the particle-number dispersion $\Delta\hat{N}^{2}$ as collective degree of freedom. These studies, which involved Skyrme, Gogny, and relativistic energy density functionals, showed that the inclusion of $\Delta\hat{N}^{2}$ on the same footing as deformation parameters strongly modifies the least-action paths, producing an overall reduction of spontaneous fission lifetimes as compared to static calculations and dynamical calculations including only mass multipole collective degrees of freedom Sadhukhan et al. (2013). Additionally, it has been demonstrated that in some nuclei pairing as a dynamic degree of freedom restores the axial symmetry at the first fission barrier Sadhukhan et al. (2014); Zhao et al. (2016). Time-dependent GCM+GOA calculation employing $\Delta\hat{N}^{2}$ as collective degree of freedom have shown that pairing correlations play an important role in shaping the fission fragment yields in induced fission Zhao et al. (2021). Also, $\Delta\hat{N}^{2}$ has been employed in beyond-mean field studies exploring the role of dynamic pairing correlations in fission Rodríguez-Guzmán et al. (2023).
The strength of pairing interactions, whether monopole, zero range, finite range, etc is usually fitted to reproduce odd-even staggering data on binding energies. The physics of staggering is far away from the physics of fission, and therefore one may wonder whether pairing strengths could be fine-tuned specifically to reproduce fission observables like spontaneous fission lifetimes. To explore this possibility, it is quite common to analyze the behavior of fission observables as a function of some parameter modulating pairing strength. In a recent example  Kouno et al. (2022, 2021), where the pairing strength in an RMF+BCS calculation was increased, a reduction of up to 2.5 MeV in the barrier height of several Pu isotopes was observed. The reduction is the consequence of a larger binding energy gain for the ground state than for the top of the barrier. This result is in contradiction with the research presented above when dynamic fission paths are characterized by stronger pairing and higher barriers at the same time. A similar result was found in Refs. Guan et al. (2021, 2023), using a macroscopic-microscopic model with BCS monopole pairing. These results were confirmed in energy density functional studies employing the Skyrme forces in the fission of 240Pu Chen et al. (2022). There, a decrease of the fission barriers up to 2.3 MeV in both symmetric and asymmetric fission channels is observed when the pairing strength is increased by 10%. The deformation of the nucleus at the scission line can also be slightly modified by modifications in the pairing interaction strength. A similar result was also discussed in Ref.  Wang et al. (2023). In microscopic Hartree-Fock-Bogoliubov (HFB) calculations with a finite range pairing force, the impact of varying the pairing strength was analyzed in Refs. Giuliani and Robledo (2013); Rodríguez-Guzmán and Robledo (2014a, b); Guzman and Robledo (2017). In those studies, increasing the pairing strength by 5%-10% can reduce the spontaneous fission lifetimes by several orders of magnitudes, particularly for those nuclei with broad and high fission barriers, mostly because of the strong damping of the collective inertias. The effect is particularly important to explain the odd-even staggering in spontaneous fission lifetimes Guzman and Robledo (2017). Moreover, when pairing is stronger, the mass parameters show smoother behavior without sharp peaks along the fission path.
All the above concerns mean field pairing correlations with their inherent spontaneous symmetry-breaking mechanism, and therefore, they do not take into account the impact of restoring particle number symmetry. Typically, symmetry restoration effects in the variation after projection (VAP) scheme also tend to increase pairing correlations Bernard et al. (2019); Rodríguez-Guzmán and Robledo (2022) but the effect is partly compensated by the reduction of pairing correlations associated with Coulomb antipairing, a component of the Coulomb interaction that is usually not considered.
@@
As the short literature overview presented here shows, the role of pairing in fission has been extensively studied in the last fifty years. The non-negligible impact of pairing in fission barriers and half-lives is currently well known. However, as various theoretical methods were used to characterize the importance of pairing correlations, results obtained from different studies cannot always be directly compared. Moreover, some contradictory conclusions were drawn from these studies, namely differences in the modification of the size of the fission barrier with pairing. These ambiguities indicate the need to systematize the knowledge regarding the impact of pairing in static fission calculations. In this work, we closely examine the three self-consistent model parameters corresponding to pairing and thoroughly investigate their impact on the description of the fission process, employing a consistent theoretical framework based on the HFB method. The paper is organized as follows: Section II briefly presents the theoretical model employed in this work. Section III presents the results of our studies. We first address the impact of the pairing gap, the particle number fluctuations, and the quenching factor parameters on fission barriers. We then discuss the results of fission half-lives for various pairing strengths and different aspects of collective inertia parameters along fission paths. In Section IV, we summarize our main conclusions. In Appendix A, we discuss the problem of the relative sign of the occupancies in the Bogoliubov transformation and its impact in the calculations present here.
@@ II Theoretical framework
The starting point for our work is the computation of potential energy surface (PES) modeling fission. We use the microscopic, self-consistent HFB model with the finite-range Gogny-type nuclear interaction. We applied the widely used and well-established D1S parametrization Gogny (1973, 1975); Dechargé and Gogny (1980); Berger et al. (1989, 1991). The computations are done in a one-center axially symmetric, deformed harmonic oscillator basis with 16 shells (including as many as 23 one-dimensional harmonic oscillator shells along the symmetry axis $n_{z\,\textup{max}}=22$). The spontaneous fission half-lives are computed using the Wentzel-Kramers-Brillouin (WKB) approximation, assuming the axially symmetric quadrupole moment $Q_{20}$ as the leading coordinate. The estimation of the probability of tunneling through the one-dimensional fission barrier requires the computation of the action $S$ along the least-energy path Brack et al. (1972):
@@
being in and out the classical turning points defined by $V(Q_{20})=E_{0}$. The potential energy appearing in the expression of the action $V(Q_{20})$ is the sum of the HFB energy and the rotational and zero point energy correction Egido and Robledo (2004); Ring and Schuck (1980). The $E_{0}$ parameter represents the quantal ground state energy obtained by solving the collective Schrödinger equation with the quadrupole moment as the degree of freedom. Although this parameter can be easily estimated by assuming a parabolic behavior of the PES around the ground state minimum and considering the collective inertia at that point, we employ a more pragmatic approach and take it equal to 0.5 MeV, which is a reasonable value. $B(Q_{20})$ is the collective inertia calculated within the perturbative cranking approximation Baran et al. (2011):
@@
where $M_{-n}$ are moments of order $n$ of the collective coordinate $Q_{20}$ (see Baran et al. (2011) for the definition). The formula for the inertia is based on the assumption that the collective operator is a one-body one, which is not true for the fluctuation on the number of particles operator $\Delta\hat{N}^{2}$ used below. To overcome this problem, the traditional approach is to use in the expression of the moments the one-body part of $\Delta\hat{N}^{2}$ in the quasiparticle basis. Although this seems a reasonable assumption, it is desirable to test it against the results obtained from a truly one-body operator, like the pairing gap one discussed below. Within the WKB approximation, the spontaneous fission half-life (in seconds) is given by Schunck and Robledo (2016):
@@
As mentioned in the Introduction I, the goal of the present work is to discuss the impact of pairing correlations on the quantities entering the action integral (2). To do so, our analysis covers the three main methods that have been traditionally employed to characterize pairing correlations strengths in fission: HFB calculations with constraints on the pairing gap parameter ($\Delta$), on the particle number fluctuation operator ($\Delta\hat{N}^{2}$), and the renormalization of the pairing interaction strength.
The pairing gap operator can be defined as $\Delta=\frac{G}{2}(\hat{P}+\hat{P}^{+})$, being $\hat{P}^{+}$ the Cooper pair creation operator, and $G=10/A$ (MeV), a parameter related to the typical strength of the monopole pairing interaction. By employing $\Delta$ as a constraining operator, it is thus possible to control the amount of pairing correlations in HFB wave functions $|\Phi\rangle$. The main advantage of employing $\Delta$ as a collective degree of freedom is its one-body nature, which allows a consistent estimation of collective inertias through Eq. (2). However, in order to employ $\Delta$ as a collective degree of freedom, specific numerical strategies must be followed in order to properly evaluate the gradient of the energy and other expectation values, as detailed in the Appendix A.
As one of the consequences of the existence of pairing correlations in a HFB calculation is the spontaneous symmetry breaking of the particle number symmetry, it is possible to employ one of the associated order parameters as a measure of pairing correlations. The most straightforward order parameter is the particle number fluctuation $\Delta\hat{N}^{2}=(\hat{N}-\langle\hat{N}\rangle)^{2}$, being $\hat{N}$ the particles number operator. This quantity is zero in the absence of pairing correlations and becomes relatively large (of the order of a few units) for strongly correlated nuclear systems. As mentioned before, a drawback related to the fluctuation in particle number is its two-body-operator character. However, owing to the gradient method used in the solutions of the HFB equation, two-body constraint operators can be handled on the same footing as one-body ones.
As a last quantity characterizing pairing correlations, we consider the quenching factor (QF), a parameter that renormalizes the strength of the pairing interaction of the Gogny force. The QF is introduced by multiplying the HFB pairing field $\Delta^{(\tau)}$ ($\tau=p,n$) by a constant factor in each iteration of the iterative solution of the HFB equation. This QF parameter is equal to 1.0 for not-modified HFB fields and takes typical values around 1.1 for enhanced pairing. It can be pictured as an analog of the pairing strength $G$ of other simple models, like BCS. In calculations employing the QF, pairing correlations can be characterized by the particle-particle energy $E_{pp}$ for protons and neutrons, which in HFB formalism is given by: